# with Knowledge Distillation for Question Answering

### v4

@@ Abstract
Multimodal multihop question answering (MMQA) requires reasoning over images and text from multiple sources. Despite advances in visual question answering, this multihop setting remains underexplored due to a lack of quality datasets. Existing methods focus on single-hop, single-modality, or short texts, limiting real-world applications like interpreting educational documents with long, multimodal content. To fill this gap, we introduce FM2DS, the first framework for creating a high-quality dataset for MMQA. Our approach consists of a 5-stage pipeline that involves acquiring relevant multimodal documents from Wikipedia, synthetically generating high-level questions and answers, and validating them through rigorous criteria to ensure data quality. We evaluate our methodology by training models on our synthesized dataset and testing on two benchmarks: MultimodalQA and WebQA. Our results demonstrate that, with an equal sample size, models trained on our synthesized data outperform those trained on human-collected data by 1.9 in exact match (EM) score on average. Additionally, we introduce M2QA-Benchwith 1k samples, the first benchmark for MMQA on long documents, generated using FM2DSand refined by human annotators. We believe our data synthesis method will serve as a strong foundation for training and evaluating MMQA models  [^1] Code and benchmark are available at: https://github.com/ServiceNow/FM2DS.
@@ 1 Introduction
Multimodal multihop question answering (MMQA) involves answering complex questions by integrating information from text, images, and tables. In real-world applications such as interpreting medical documents, this challenge is amplified by the need to reason over long, multimodal content. Current methodologies in MMQA typically leverage in-context learning methods, prompting LVLMs to retrieve relevant information from multimodal sources (Tejaswi et al., 2024) and then perform reasoning (Yang et al., 2023). However, these models often demand significant computational resources due to their large parameter counts, making them costly to deploy even during inference (Ye et al., 2024). This limitation emphasizes the need for more efficient frameworks that can operate effectively with minimal annotated data. A practical solution is to use a smaller model capable of both retrieving the necessary information from sources and performing reasoning. This can be achieved by fine-tuning the model on a MMQA dataset. Existing datasets often rely on short snippets or repetitive templates, limiting generalizability to complex settings with long texts and multiple modalities (Chang et al., 2021; Talmor et al., 2021; Jiang et al., 2024; Chen et al., 2024a). Additionally, creating new similar datasets is challenging, requiring extensive human annotation (Lu et al., 2022b; Chen et al., 2023a).
@@
In this work, we propose FM2DS, a novel data synthesis framework designed specifically for MMQA over long documents. Our approach synthesizes MMQA data from documents that are interconnected through various relationships, such as thematic similarities or sequential events. This framework leverages naturally occurring document relationships and requires minimal hand-crafted data, thereby broadening the range of reasoning types used in question generation. As illustrated in Figure 1, FM2DS enables the generation of non-templated question-answer pairs based on full documents rather than brief information snippets. The data generated by our method FM2DS includes query component— a step-by-step guide for retrieving relevant information from multiple documents — enabling smaller models trained on this synthesized data to learn how to tackle complex questions in a manner similar to larger models. This methodology allows users to create a custom MMQA dataset with fewer than ten human-annotated samples, thereby facilitating the fine-tuning of smaller LVLMs for specific applications.
@@
FM2DS leverages Wikipedia’s extensive knowledge base and hyperlink structure, to select document pairs with shared topical relevance or hyperlink connections and prompt LVLMs to perform question generation, question answering, and query generation. We incorporate validation steps to enhance the quality of the generated data and discard any outputs that are factually incorrect. Through empirical evaluation on established MMQA benchmarks, we show that FM2DS significantly improves model performance, achieving on average a 1.9 exact match (EM) score improvement across two benchmarks: MultimodalQA and WebQA. Our key contributions are: (I) introducing a new framework for synthesizing high-quality MMQA training data for VLMs; (II) using a robust validation pipeline to ensure data quality; (III) introducing a challenging MMQA benchmark requiring reasoning over multiple modalities and sources; and (IV) showing that models fine-tuned on our synthetic data outperform those trained on human-labeled datasets, advancing MMQA while reducing manual effort.
@@ 2 Related Work
Figure 2: The five-stage pipeline for FM2DS begins with retrieving related Wikipedia documents based on hyperlinks and topics (Stage 1). Next, we select few-shot examples from MultiModalQA Talmor et al. (2021) (Stage 2). Stage 3 generates and validates questions to ensure they are answerable, multihop, and multimodal. In Stage 4, answers are generated and validated. Finally, Stage 5 produces and validates queries to guide complex MMQA.
@@ 3 Proposed Method: FM2DS
Our five-stage pipeline for FM2DS (Figure 2) synthesizes high-quality multimodal QA pairs. It begins by grouping documents via topic matching and Wikipedia hyperlinks, followed by few-shot sample selection, question synthesis, answer generation, and query construction, each with their built-in validation. See Appendix I for examples.
@@ 3.1 Stage 1: Creating a Pool of Related Documents
We collect relevant documents from Wikipedia using the WikiWeb2M dataset (Burns et al., 2023), which includes nearly 2 million pages. Documents are linked via two methods: hyperlinks and latent topics identified through multimodal topic modeling with the Multimodal-Contrast model (González-Pizarro & Carenini, 2024). Since Multimodal-Contrast can’t handle long documents, we split each document into shorter segments containing at most one image, applied topic modeling to each segment, then merged the results and removed duplicates. This combination captures both clear and subtle relationships across documents, integrating textual and visual information.
@@ 3.2 Stage 2: Creating Few-Shot Samples
We sample multihop questions from MultiModalQA dataset (Talmor et al., 2021), which requires reasoning across text, images, and tables. As our samples are based on full documents rather than short information snippets like in MultimodalQA, we crawled the complete Wikipedia HTML pages using the entity links provided in MultimodalQA, which are associated with the dataset’s examples. We then compiled few-shot samples using these full HTML pages—complete with images and tables—paired with their corresponding questions. We randomly select up to three samples for question generation in our experiments.
@@ 3.3 Stage 3: Question Generation and Validation
We use GPT-4-turbo (OpenAI et al., 2024) to generate multihop, multimodal questions from few-shot samples based on MultiModalQA few-shot examples. Due to context limitations, inputs are limited to grouped sets containing 2 or 3 documents. Our prompt (Appendix A) is designed to ensure that questions require reasoning across all documents and at least two modalities, avoiding unrelated combinations such as: “How did Einstein contribute to relativity and when was Princeton established?”—a question that spans multiple documents but lacks meaningful multihop reasoning.
Question Type
In what year did Mike Tyson become the youngest heavyweight champion, and who is the president of the United States? Unrelated Facts
Question Type
In what year did Mike Tyson become the youngest heavyweight champion, and who was the president of the United States at that time? Related Facts, Open-ended
Question Type
Who was the president of the United States when Mike Tyson became the youngest heavyweight champion? Concise Multihop Question
@@
We used LLama-3.1-8B (Abhimanyu Dubey et al., 2024) to decompose questions and check if parts could be answered with a single document. If all parts of the question were answerable with a single document, we discarded such question that include unrelated facts (see Unrelated Facts example in Table 1). Alternatively, we retained only the initial component that required information from multiple documents to ensure the revised question met the multihop criteria. However, a potential issue was that, even when the facts were related, the questions could still become open-ended, requiring explanations or combined answers (see example Related Facts, Open-ended in Table 1). In order to follow the standard of question answering, and make the evaluation process easier, we used GPT-4o to rephrase the question without conjunctions while maintaining its multihop nature, resulting in Concise Multihop Question (Table 1).
@@ 3.4 Stage 4: Answer Generation and Validation
We validated answers using named entity recognition (NER) and relation extraction, following prior work (Rajpurkar et al., 2018; Fabbri et al., 2022). NER ensured key entities and numbers in the answer matched the documents, while relation extraction verified that entity relationships were consistent with the source content (via Spacy (Wu & He, 2019)). For including image content, we used the same question-related caption generated by GPT-4o (e.g., noting a building’s color if relevant to the question) similar to answer generation. To reduce hallucinations, we prompted GPT-4o five times and accepted answers only if all outputs (5/5) agreed. To evaluate the effectiveness of our answer validation process, we conducted a human study to assess the quality of the filtered questions and answers. The results of this evaluation are presented in Section 6.
@@ 3.5 Stage 5: Query Generation and Validation
We generate queries using GPT-4o based on the question-answer pairs and related documents to enhance retrieval effectiveness. These queries help focus the search on specific, relevant information needed to answer the questions accurately. By narrowing down the content, we can extract key details such as named entities, relationships, and contextual cues aligned with the question. This targeted approach ensures that the generated answers are not only concise and accurate but also directly grounded in evidence from the documents.
@@
4 Proposed Benchmark: M2QA-Bench
We introduce M2QA-Bench, a benchmark of 1k diverse Q&A pairs to evaluate LVLMs on complex MMQA with full documents. Unlike templated datasets (Talmor et al., 2021), questions are varied and challenging (see Appendix G for details on complexity and diversity). Answering requires cross-modal reasoning and information extraction from full documents, including images and tables. See Table 2 for key statistics (more in Appendix G) and Appendix I for samples generated by FM2DS for M2QA-Bench. To create this benchmark, we used the FM2DS pipeline to generate 1,200 samples, which were verified by three annotators for correctness, multihop reasoning, multimodality, and answer accuracy. Each sample was scored 1 (valid) or 0 (invalid). This annotation required minimal human effort (2.2 min/question on average) due to structured queries. Samples averaging below 0.75 were removed, leaving 1,142 (i.e removing only 5% of the total); we then randomly selected 1,000 for the benchmark to ensure consistency in evaluation and reduce potential sampling bias. Annotator agreement (Fleiss’ Kappa (Fleiss, 1971)) was 0.83.
@@ 5 Experiments and Results
This section compares our synthesized dataset to human-annotated ones. All experiments used one in-context example during synthesis (see Appendix C for effects of varying number in-context examples). GPT-4o was used in the pipeline (Appendix D shows results with other LVLMs). Models were evaluated using Exact Match (EM) for accuracy and F1 for partial match quality. Further experimental details can be found in Appendix B.
@@
Table 3: Comparison of synthetic and human-annotated data across various models. Smaller models were evaluated with 5k, 10k, and full training sets (23.8k for MultiModalQA, 34.2k for WebQA), while larger models used 10k and full sets. For each model, the listed synthetic sample size is the smallest (divisible by 1k) that outperforms the same model trained on the full human-labeled set. For real samples, we used the WebQA training set for testing on the WebQA test set, and similarly for MultiModalQA. “None” indicates default pretrained models. For more results, see Appendix E.

### v5

@@ Abstract
Multimodal multihop question answering (MMQA) requires reasoning over images and text from multiple sources, an essential task for many real-world applications. Despite advances in visual question answering, the multihop setting remains underexplored due to the lack of quality datasets. Existing methods focus on single-hop, single-modality, or short texts, limiting real-world applications like interpreting educational documents with long, multimodal content. To fill this gap, we introduce FM2DS, the first framework for creating a high-quality dataset for MMQA. Our approach consists of a five-stage pipeline that involves acquiring relevant multimodal documents from Wikipedia, synthetically generating high-level questions and answers, and validating them through rigorous criteria to ensure data quality. We evaluate our methodology by training models on our synthesized dataset and testing on two benchmarks: MultimodalQA and WebQA. Our results demonstrate that, with an equal sample size, models trained on our synthesized data outperform those trained on human-collected data by 1.9 in exact match (EM) score on average. Additionally, we introduce M2QA-Bench with 1k samples, the first benchmark for MMQA on long documents, generated using FM2DS  and refined by human annotators [^1] Code is publicly available at: https://github.com/ServiceNow/FM2DS..
[Uncaptioned image]FM2DS: Few-Shot Multimodal Multihop Data Synthesis
with Knowledge Distillation for Question Answering
Amirhossein Abaskohi1,2, Spandana Gella2, Giuseppe Carenini1, Issam H. Laradji1,2 1 Department of Computer Science The University of British Columbia V6T 1Z4, Vancouver, BC, Canada 2 ServiceNow Research
@@ 1 Introduction
Multimodal multihop question answering (MMQA) involves answering complex questions by integrating information from text, images, and tables. In real-world applications such as interpreting medical documents, this challenge is amplified by the need to reason over long, multimodal content. Current methodologies in MMQA typically leverage in-context learning methods, prompting large vision language models (LVLMs) to retrieve relevant information from multimodal sources (Tejaswi et al., 2024) and then perform reasoning (Yang et al., 2023). However, these models often demand significant computational resources due to their large parameter counts, making them costly to deploy even during inference (Ye et al., 2024). This limitation emphasizes the need for more efficient frameworks that can operate effectively with minimal annotated data. A practical solution is to use a smaller model capable of both retrieving the necessary information from sources and performing reasoning. This can be achieved by fine-tuning the model on a MMQA dataset. Fine-tuning enables domain specialization, allowing the model to adapt to specific areas of interest. It also requires significantly less compute and memory compared to large commercial models. Finally, this approach reduces privacy concerns by enabling training on sensitive domains such as legal, medical, or proprietary data that models like GPT-4o OpenAI et al. (2024) cannot access. Existing datasets often rely on short snippets or repetitive templates, limiting generalizability to complex settings with long texts and multiple modalities (Chang et al., 2021; Talmor et al., 2021; Jiang et al., 2024; Chen et al., 2024a). Additionally, creating new similar datasets is challenging, requiring extensive human annotation (Lu et al., 2022a; Chen et al., 2023a).
@@
In this work, we propose FM2DS, a novel data synthesis framework designed specifically for MMQA over long documents. Our approach synthesizes MMQA data from documents that are interconnected through various relationships, such as thematic similarities or sequential events. This framework leverages naturally occurring document relationships and requires minimal hand-crafted data, thereby broadening the range of reasoning types used in question generation.
As illustrated in Figure 1, FM2DS enables the generation of non-templated question-answer pairs based on full documents rather than brief information snippets. The data generated by our method FM2DS includes query component - a step-by-step guide for retrieving relevant information from multiple documents - enabling smaller models trained on this synthesized data to learn how to tackle complex questions in a manner similar to larger models. This methodology allows users to create a custom MMQA dataset with fewer than ten human-annotated samples, thereby facilitating the fine-tuning of smaller LVLMs for specific applications.
@@
FM2DS leverages Wikipedia’s extensive knowledge base and hyperlink structure to select document pairs with shared topical relevance or hyperlink connections and prompt LVLMs to perform question generation, question answering, and query generation. We incorporate validation steps to enhance the quality of the generated data and discard any outputs that are factually incorrect. Through empirical evaluation on established MMQA benchmarks, we show that FM2DS significantly improves model performance, achieving on average a 1.9 exact match (EM) score improvement across two benchmarks: MultimodalQA and WebQA.
Our key contributions are: (I) introducing a new framework for synthesizing high-quality MMQA training data for LVLMs; (II) using a robust validation pipeline to ensure data quality; (III) introducing a challenging MMQA benchmark requiring reasoning over multiple modalities and sources; and (IV) showing that models fine-tuned on our synthetic data outperform those trained on human-labeled datasets, advancing MMQA while reducing manual effort.
@@ 2 Related Work
Figure 2: The FM2DS pipeline consists of five stages for generating high-quality multihop multimodal QA samples. In Stage 1, a pool of related Wikipedia documents is retrieved by leveraging topic similarity and hyperlink connections to ensure contextual richness. Stage 2 selects few-shot in-context examples from the MultiModalQA dataset Talmor et al. (2021) to guide generation. Stage 3 focuses on question generation (3.1) and validation (3.2), ensuring questions require multihop reasoning, are answerable, and grounded in both text and images. Stage 4 generates (4.1) and validates (4.2) answers through named entity alignment, relation consistency, and hallucination checks. Finally, Stage 5 generates (5.1) and validates (5.2) retrieval queries to collect diverse and relevant supporting documents. The resulting samples are saved in a structured format for use in MMQA training and evaluation.
@@ 3 Proposed Method: FM2DS
Our five-stage pipeline for FM2DS (Figure 2) synthesizes high-quality multimodal QA pairs. It begins by grouping documents via topic matching and Wikipedia hyperlinks, followed by few-shot sample selection, question synthesis, answer generation, and query construction, each with their built-in validation. See Appendix K for examples.
@@ 3.1 Stage 1: Creating a Pool of Related Documents
We collect relevant documents from Wikipedia using the WikiWeb2M dataset (Burns et al., 2023), which includes nearly 2 million pages. Documents are linked via two methods: hyperlinks and latent topics identified through multimodal topic modeling with the Multimodal-Contrast model (González-Pizarro and Carenini, 2024). Since Multimodal-Contrast can not handle long documents, we split each document into shorter segments containing at most one image, apply topic modeling to each segment, then merge the results and remove duplicates. This combination captures both clear and subtle relationships across documents, integrating textual and visual information.
[…]
@@ 3.2 Stage 2: Creating Few-Shot Samples
We sample multihop questions from the MultiModalQA dataset (Talmor et al., 2021), which requires reasoning across text, images, and tables. As our samples are based on full documents rather than short information snippets like in MultimodalQA, we crawled the complete Wikipedia HTML pages using the entity links provided in MultimodalQA, which are associated with the dataset’s examples. We then compiled few-shot samples using these full HTML pages, complete with images and tables, paired with their corresponding questions. We randomly select up to three samples for question generation in our experiments.
@@ 3.3 Stage 3: Question Generation and Validation
We use GPT-4-turbo (OpenAI et al., 2024) to generate multihop, multimodal questions from few-shot samples based on MultiModalQA few-shot examples. Due to context limitations, inputs are limited to grouped sets containing 2 or 3 documents. Our prompt (see Appendix A) is designed to ensure that questions require reasoning across all documents and at least two modalities, avoiding unrelated combinations such as: “How did Einstein contribute to relativity and when was Princeton established?"; a question that spans multiple documents, but lacks meaningful multihop reasoning.
@@
We used LLama-3.1-8B (Abhimanyu Dubey et al., 2024) to decompose questions and check if parts could be answered with a single document. If all parts of the question were answerable with a single document, we discarded such question that include unrelated facts (see Unrelated Facts example in Table 1). Otherwise, we retained only the the parts of the questions that required information from multiple documents to ensure the revised question met the multihop criteria. However, a potential issue was that, even when the facts were related, the questions could still become open-ended, requiring explanations or combined answers (see example Related Facts, Open-ended in Table 1). In order to follow the standard of question answering, and make the evaluation process easier, we used GPT-4o to rephrase the question without conjunctions while maintaining its multihop nature, resulting in Concise Multihop Question (Table 1).
@@ 3.4 Stage 4: Answer Generation and Validation
We validated answers using named entity recognition (NER) and relation extraction, following prior work (Rajpurkar et al., 2018; Fabbri et al., 2022). NER ensured key entities and numbers in the answer matched the documents, while relation extraction verified that entity relationships were consistent with the source content (via Spacy (Wu and He, 2019)). For including image content, we used the same question-related caption generated by GPT-4o (e.g., noting a building’s color if relevant to the question) similar to answer generation. To reduce hallucinations, we prompted GPT-4o five times and accepted answers only if all outputs (5/5) agreed. To evaluate the effectiveness of our answer validation process, we conducted a human study to assess the quality of the filtered questions and answers. The results of this evaluation are presented in Section 6.
@@ 3.5 Stage 5: Query Generation and Validation
We generate queries using GPT-4o based on the question-answer pairs and related documents to enhance retrieval effectiveness. These queries guide the smaller model trained on FM2DS-generated data to retrieve specific and relevant information, improving its ability to answer questions accurately. By narrowing down the content, we can extract key details such as named entities, relationships, and contextual cues aligned with the question. This targeted approach ensures that the generated answers are not only concise and accurate, but also directly grounded in evidence from the documents.
@@
4 Proposed Benchmark: M2QA-Bench
We introduce M2QA-Bench, a benchmark of 1k diverse Q&A pairs to evaluate LVLMs on complex MMQA with full documents. Unlike templated datasets (Talmor et al., 2021), questions are varied and challenging (see Appendix I and Appendix F for details on diversity and complexity). Answering requires cross-modal reasoning and information extraction from full documents, including images and tables. See Table 2 for key statistics (more in Appendix I) and Appendix K for samples generated by FM2DS for M2QA-Bench. To create this benchmark, we used the FM2DS pipeline to generate 1,200 samples, which were verified by three annotators for correctness, multihop reasoning, multimodality, and answer accuracy. Each sample was scored 1 (valid) or 0 (invalid). This annotation required minimal human effort (2.2 min/question on average) due to structured queries. Samples averaging below 0.75 were removed, leaving 1,142 (i.e removing only 5% of the total); we then randomly selected 1,000 for the benchmark to ensure consistency in evaluation and reduce potential sampling bias. Annotator agreement (Fleiss’ Kappa (Fleiss, 1971)) was 0.83.
@@ 5 Experiments and Results
This section compares our synthesized dataset to human-annotated ones. All experiments used one in-context example during synthesis (see Appendix D for effects of varying number in-context examples). GPT-4o was used in the pipeline (Appendix E shows results with other LVLMs). Models were evaluated using Exact Match (EM) for accuracy and F1 for partial match quality. Further experimental details can be found in Appendix B.
5.1 Details of Synthesized Experimental Training Data
We synthesize an 18k-sample training set with FM2DS under the 3-shot setting, using a 20-example few-shot pool (10 from MultimodalQA, 10 from WebQA) to guide style and difficulty. Table 3 reports aggregate corpus statistics. The dataset is inherently multimodal: the majority of questions reference images, tables, or both. In most of our experiments, we use either a 5k or 10k subset of this training data to balance efficiency and performance. Larger subsets (>10>10k) are employed only in special cases where we aim to identify the minimum amount of synthesized data required to surpass training on the full ground-truth training set of the respective test benchmark. This threshold varies depending on the model and the evaluation dataset. To ensure training utility, we enforce basic validity checks during synthesis (e.g., modality availability, answerability, and cross-source consistency) and remove low-confidence or duplicate generations. These design choices yield a balanced, compact corpus that emphasizes multimodal reasoning without sacrificing clarity. See Appendix C for data generation cost details.
Statistic Value
Image Modality Percentage 64.55%
Table Modality Percentage 64.4%
Both Image and Table Modality Percentage 42.2%
Average Question Length (Word) 22.5
Average Answer Length (Word) 1.94
Average Source Documents Per Question 2.1
Table 3: Key statistics of the 18k synthesized training samples generated by FM2DS under the 3-shot setting. While we generated 18k samples in total, most experiments use 5k or 10k subsets, with larger subsets employed only when testing the minimum required size to outperform training on the original ground-truth training data of the respective benchmarks.
5.2 Training Details
Structured Query Format for Knowledge Distillation.
To promote explicit and grounded multimodal reasoning, we train models in a structured query–answer format where each prediction requires the model to (I) identify the relevant modality (image, table, or both), (II) extract or cite supporting evidence, and (III) generate the final answer. Figures 10 and 11 in Appendix K illustrate examples of queries. The model is supervised to generate both intermediate queries and final answers using a masked language modeling (MLM) loss, preventing shortcut learning and ensuring that reasoning is explicitly tied to content.
We align this framework with knowledge distillation from a stronger teacher model (GPT-4o). The teacher produces both structured queries (chain-of-thought style reasoning) and answers, which serve as distillation targets. A validation step filters out hallucinated or factually inconsistent teacher queries before use. The student model is then optimized to mimic the teacher’s reasoning and answer trajectories, similar in spirit to reasoning-focused distillation methods such as DeepSeek-R1-Distill DeepSeek-AI et al. (2025), but adapted for multimodal multihop QA.
Queries are essential for guiding models to use the provided context, as without explicit multimodal grounding they struggle to answer reliably despite strong pretrained knowledge. As demonstrated in Appendix F, removing supporting context causes sharp performance drops across all baselines, highlighting the importance of grounding multimodal multihop question answering.
Training Objective.
Given an input $(x,y,r)$ where $x$ is the multimodal context (including the question and documents), $r$ is the structured query (teacher reasoning), and $y$ is the ground-truth answer, the model is optimized with a joint objective:
$\mathcal{L}=\mathcal{L}_{\text{CLM}}(r|x)+\mathcal{L}_{\text{CLM}}(y|x,r),$
where $\mathcal{L}_{\text{CLM}}$ denotes the causal language modeling loss. The first term enforces generation of factually grounded queries, while the second supervises answer generation conditioned on both the input and queries. In the distillation setting, teacher outputs $(r^{*},y^{*})$ replace $(r,y)$ to guide the student model toward teacher-quality reasoning and answering.
@@ 5.2 Training Details
Claude-4-Sonnet None 78.25 82.92 None 82.18 87.61
Table 4: Comparison of synthetic and human-annotated data across various models. We generated 18k synthetic training samples in total, but models are typically trained on 5k or 10k subsets for efficiency. Larger subsets (>10>10k) are only used in cases where we seek the minimum number of synthetic samples needed to outperform training on the full human-labeled set. For smaller models, we evaluate with 5k, 10k, and full training sets (23.8k for MultiModalQA, 34.2k for WebQA), while larger models use 10k and full sets. For each model, the listed synthetic sample size is the smallest (divisible by 1k) that surpasses the same model trained on the full human-labeled set. For real samples, we used the WebQA training set for testing on the WebQA test set, and similarly for MultiModalQA. “None” indicates default pretrained models. For extended results, see Appendix G.